OpenStudy (ggrree):

Find the are of the region enclosed between the graphs y = 3x and y =3x^2 - 6x I know the answer, and I think I know how to solve it, but I'm having trouble computing the actual answer. Help would be appreciated.

OpenStudy (anonymous):

= 13.5

OpenStudy (nikvist):

\[y=3x\quad,\quad y=3x^2-6x\]\[3x^2-6x=3x\]\[3x^2-9x=0\quad;\quad x_1=0\quad,\quad x_2=3\]\[S=\int\limits_0^3(9x-3x^2)dx\]

OpenStudy (ggrree):

is that all there is to it, nikvist ? A portion of the region lies below the x axis, so I think that your answer is incorrect.

OpenStudy (ggrree):

that's correct, chlorophyll! Could you please show your work? I am having trouble getting there.

OpenStudy (nikvist):

check my work again. I think my work is correct

OpenStudy (anonymous):

Just exactly as the nikvist's progress!

OpenStudy (anonymous):

Take derivative then plug x = 3 in!

OpenStudy (anonymous):

In mean Integral!

OpenStudy (ggrree):

I am not understanding how he got \[\int\limits_{0}^{3} 9x - 3x^2 dx\]. Why are the signs reversed here? shouldn't it at the very least be \[\int\limits_{0}^{3} 3x^2 - 9x dx\] ? since that is what he wrote the line before.

OpenStudy (ggrree):

I computed and got a negative value, which doesn't make sense :(

OpenStudy (anonymous):

since when you sketch the graph, y = 3x is above y = 3x^2 - 6x

OpenStudy (anonymous):

That's why you subtract: 3x - ( 3x^2 -6x)

OpenStudy (ggrree):